OpenStudy (anonymous):

SOS!!!!! One number is 5 more than twice another and their sum is 23. Which of the following system of equations represents the word problem? A) x + y = 23 and y = 2x + 5 B) y = x + 5 and y + 2(x + 5) = 23 C)y = 2(x + 5) and x + y = 23

OpenStudy (anonymous):

B0

OpenStudy (anonymous):

rory gtfo

OpenStudy (anonymous):

fight me irl

OpenStudy (anonymous):

rsad

OpenStudy (imstuck):

What is going on? Do you need help with this or what?

OpenStudy (anonymous):

shes fine

OpenStudy (anonymous):

@IMStuck Please!!

OpenStudy (anonymous):

@IMStuck Don't listen to him, I need help!

OpenStudy (imstuck):

You need help, I will help, ok? Let's call the first unknown in the first set of "instructions" y. So they say that y is 5 more than twice another number. Let's call the other number x. So far then we have, in words, y = 5 + 2x. The word "is" translates to an = sign, and "twice more" is 2 times, and "5 more" is 5 +. Do you get that so far? The first equation is y = 5 + 2x

OpenStudy (imstuck):

Now you have identified your unknowns as x and y, and it tells you that the sum of them is 23. How would you translate that into an equation?

OpenStudy (anonymous):

would it be y = 5 + 2x = 23?

OpenStudy (anonymous):

no thats wrong

OpenStudy (imstuck):

Actually, the sum of means adding, and if you need the sum of x and y and it is equal to 23 you have x + y = 23

OpenStudy (anonymous):

Oh, that makes sense!

OpenStudy (imstuck):

Yeah. And of course the first equation, y = 5 + 2x can be rearranged to say y = 2x + 5. That doesn't change anything when you write it like that. so do you see those two equations in a choice of answers up there?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Thank you!

OpenStudy (imstuck):

Cool! Also, if I were you, I would block that person. I do not appreciate seeing that foul garbage written here while helping you solve math problems. Just saying! Anytime you need help, just ask! : )

OpenStudy (anonymous):

Thank you :)